OpenStudy (anonymous):

Malia is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Malia's inputs and the corresponding outputs displayed by the calculator: (1, 3), (2, 7), (3, 11), (4, 15) Which of the following functions best represents the rule that the calculator uses to display the outputs? f(n) = 5n - 1 f(n) = 5n + 1 f(n) = 4n + 1 f(n) = 4n - 1

OpenStudy (anonymous):

@SolomonZelman , @welshfella

OpenStudy (welshfella):

plug in the given values into each option and see which one fits

OpenStudy (anonymous):

I found that when I did that I got D to be y answer since they all matched.

OpenStudy (welshfella):

for example first choice 5n - 1 input 1 output 3 input n = 1 output 5n - 1 = 5(1) - 1 = 4 but it should be 3 so its not this option

OpenStudy (welshfella):

thats right its D

OpenStudy (anonymous):

Thanks!

OpenStudy (welshfella):

yw